# On the Resilience of Multi-Agent Systems with Malicious Agents

### v2

@@ 4.3 RQ2: Impact of Downstream Tasks
Tasks requiring rigor and formalization, such as code generation and math, are more sensitive to agent errors and exhibit lower resilience compared to translation and text evaluation. Code generation and math demand greater objectivity than the more subjective tasks of translation and text evaluation. Fig. 3b illustrates the impact of AutoTransform and AutoInject across different downstream tasks. We also present the performance of single-agent using GPT-3.5 with the prompts listed in §B.2, for a clearer comparison. The results indicate several conclusions: (1) Multi-agent systems can outperform single-agent settings, but their performance may decline to similar or worse levels when affected by malicious agents. (2) Objective tasks benefit more from multi-agent collaboration, while subjective tasks gain less. Additionally, errors in subjective tasks are often overlooked by other agents due to the lack of rigorous correctness standards. (3) In terms of system resilience, tasks ranked from least to most vulnerable are: code generation, math, translation, and text evaluation. Even minor errors in the first two tasks, particularly in code generation, significantly affect rigor and formalization. Conversely, the latter two tasks are less sensitive to minor variations in a single agent’s output. (4) AutoTransform and AutoInject perform similarly across most tasks, except in code generation.
@@
Injecting errors can surprisingly improve performance on downstream tasks. We find that certain multi-agent collaboration systems, such as MAD, Camel, and AgentVerse, benefit from deliberately injected errors rather than being hindered by them. Fig. 4 shows the performance changes of MAD with AutoInject. Additionally, Camel’s text evaluation performance increases from $43.8\%$ to $49.5\%$, while AgentVerse’s translation performance also rises from $43.8\%$ to $49.5\%$.
Figure 4: An increase of accuracies observed on MAD against AutoInject on each task.
@@
We now present two scenarios where deliberately introduced errors enhance system performance. (1) Double Checking: Introducing an obvious error prompts the system (i.e., other agents) to require the malicious agent to produce another message to correct the erroneous code. This process not only corrects the injected error but also fixes pre-existing errors in the original code, thereby increasing the likelihood of task completion. (2) Divergent Thinking: Systems like MAD, which incorporate a debate mechanism, may sometimes get trapped in repetitive loops due to relying on the same LLMs as their backbone, resulting in stagnant discussions. By intentionally adding significant errors that shift the original distribution, we can help agents break free from these limitations. This finding aligns with and extends the conclusions from Du et al. (2024) and Liang et al. (2024) that agents with diverse opinions can facilitate problem solving. Additionally, this mechanism explains why AutoInject can improves performance, while AutoTransform, which lets agents produce errors themselves, cannot.
@@ 4.4 RQ3: Impact of Error Rates
Increasing the number of erroneous messages causes a larger performance drop than the number of errors within a message. Since AutoTransform lacks precise control over error rates and types, we focus on AutoInject for RQ3 and RQ4. Fig. 5a presents two experiments: one with a fixed $P_{e}=0.2$ and varying $P_{m}$ at $0.2$, $0.4$, and $0.6$, labeled “Erroneous Message;” The other with a fixed $P_{m}=0.2$ and varying $P_{e}$ at $0.2$, $0.4$, and $0.6$, labeled “Errors per Message.” For “Erroneous Message,” as $P_{m}$ increases, the task performance consistently decreases. Regarding the error ratio in a single message: (1) In contrast to $P_{m}$, the performance reached a bottleneck as $P_{e}$ increases from $0.4$ to $0.6$. (2) As $P_{e}$ increases, performance decreases, implying that while higher error rates make errors more noticeable, the agent system struggles to correct the increasing number of errors. An exception is observed when increasing $P_{e}$ from $0.4$ to $0.6$, resulting in a performance increase in three systems (MetaGPT, Self-collab, MAD). This occurs because excessive errors in a single message become noticeable, prompting other agents to request corrections. This phenomenon highlights the importance of stealth in introducing errors.
(a) Error rate in AutoInject.
(b) Error type in AutoInject.
Figure 5: Performance drops of the six multi-agent systems on selected downstream tasks.
@@ 4.5 RQ4: Impact of Error Types
Semantic errors cause a greater performance drop than syntactic errors. Fig. 5b presents the performance decline caused by syntactic and semantic errors across six systems, including the average. Most systems handle syntactic errors more effectively than semantic errors. This likely stems from LLMs excelling at identifying syntactic errors due to their extensive training on code corpora, where such errors differ from the training data distribution. In contrast, semantic errors resemble correct code in distribution, requiring a deeper task understanding (e.g., whether the loop should start at 1 or 0) for accurate identification. For instance, in the Camel system, syntax errors in the Assistant agent prompt the User agent to instruct “correct the mistakes in the code,” forcing the Assistant agent to rectify the code. Notably, syntactic errors have minimal impact on Self-collab and MAD; in fact, MAD shows improved performance with injected syntactic errors. Self-collab utilizes an external compiler to ensure code execution, while MAD employs a higher-level agent (the Judge agent) to produce the final result.
4.6 Case Study
Current LLMs prioritize natural language over code. Fig. 6b illustrates that distraction comments can mislead LLMs into accepting incorrect code as correct across all six systems studied. This indicates that the systems tend to prioritize comments over the actual code. In the example, the system detects an error in the code when no comments are present. However, when a comment stating “the bug had been corrected” is added, the system overlooks the error and proceeds with the next task. AutoTransform exploits this characteristic of LLMs to execute successful attacks.
@@ 4.6 Case Study
4.7 Other Factors
Impact of Malicious Roles
Previous experiments in §4 focus on polluting the agents directly responsible for the work, rather than those who delegate tasks to other agents. To examine the impact of polluting different types of agents and the generalizability of our AutoTransform on agents with varying roles, this section investigates the effects of polluting high-level agents. Specifically, we apply AutoTransform to the User and Assistant agents in Camel, and the Product Manager and Engineer agents in MetaGPT. The results of these systems completing code generation tasks are shown in Fig. 7a. The conclusions are as follows: (1) AutoTransform is applicable to agents with different profiles or functionalities, effectively disrupting collaboration. (2) Polluting higher-level task distributors results in a greater performance drop for both systems. The second finding aligns with our intuition that instructors controlling the broader aspects are more crucial. For example, in Camel, the Assistant agent struggles to recognize “toxic” instructions from the User agent due to its role of merely following instructions.
(a) The accuracy of polluting different roles.
(b) Average rounds for Camel on HumanEval.
Figure 7: The two factors studied in §4.7. (a) Impact of applying AutoTransform on different roles in MetaGPT and Camel. (b) Correlation of average rounds with the correctness of code.
Impact of Numbers of Rounds
Another intuition is that increased agent involvement (i.e., more rounds) enhances system resilience. To eliminate the influence of additional agents, we focus on Camel which has only two agents who take turn to speak. We compute the average number of rounds for both correct and incorrect code generation. As shown in Fig. 7b, without injected errors, the average rounds for code passing HumanEval is $9.31$, while for non-passing code, it is $9.79$. After injecting errors with AutoInject, these averages change to $8.89$ and $11.57$, respectively. This suggests that error injection leads the system to complete easier examples with shorter conversations, while spending more time on harder cases without improvement. However, this contradicts the intuition that the number of rounds may correlate with system resilience, aligning with the finding that the effect of the number of agents or rounds is limited Amayuelas et al. (2024).
@@ 5 Improving System Resilience
Based on our experimental observations and findings, we propose two strategies for improving resilience in multi-agent collaboration systems, defending against malicious agents.
Defense Methods
The core idea behind our defense methods involves adding a correction mechanism within the system. We explore two variants, the “Challenger” and the “Inspector.” The “Challenger,” akin to our AutoTransform, is an additional description of functionalities added in agent profiles. This method addresses the limitation that many agents can only execute assigned tasks and may not address certain problems they encounter, although they usually have the knowledge to. By empowering agents to challenge the results of others, we enhance their problem-solving capabilities. This is because most current multi-agent systems use the same LLM as the backbone for all agents, indicating their underlying ability to partially solve tasks outside their specialization.
In contrast, the “Inspector,” similar to our AutoInject, is an additional agent that intercepts all messages spread among agents, checks for errors, and corrects them. This method draws inspiration from the “Police” agent in Zhang et al. (2024). Detailed prompts for the “Challenger” and “Inspector” methods can be found in §B.5 and §B.6, respectively, in the appendix.
Figure 8: Performance of defense methods, “Challenger” and “Inspector,” on code generation task.
@@ Results
We apply these defense strategies to the two weaker architectures: the linear (Self-collab) and the flat (Camel). Results for the vanilla model, error injection with AutoInject, and the two defense methods are shown in Fig. 8. Both defense methods improve performance against AutoInject, though they do not restore it to the original level. With Challenger, we recover $87.93\%$ of the performance loss caused by malicious agents (improving from $40.85\%$ to $71.95\%$, compared to the original $76.22\%$). However, no definitive conclusion can be drawn regarding which method is superior, as Inspector outperforms Challenger on Camel. We recommend testing both methods in practice.
@@ 6.1 Multi-Agent Systems
LLMs enhance multi-agent systems through their exceptional capability for role-play Wang et al. (2024b). Despite utilizing a same architecture, like GPT-3.5, distinct tasks benefit from tailored in-context role-playing prompts Min et al. (2022). Besides the six frameworks selected in this study, researchers have been exploring multi-agent collaboration in downstream tasks or simulated communities. ChatEval Chan et al. (2024) is a multi-agent debate system for evaluating LLM-generated text, providing a human-like evaluation process. ChatDev Qian et al. (2024) uses a linear structure of several roles to address code generation tasks. AutoGen Wu et al. (2023) offers a generic framework for building diverse applications with multiple LLM agents. AutoAgents Chen et al. (2023) enables dynamic generation of agents’ profiles and cooperation, evaluated on open-ended QA and creative writing tasks. Zhou et al. (2023) support planning, memory, tool usage, multi-agent communication, and fine-grained symbolic control for multi-agent or human-agent collaboration. Additionally, there are studies simulating daily life or conversations Park et al. (2023); Zhou et al. (2024b), and multi-agent competition Huang et al. (2024); Liu et al. (2024); Liang et al. (2023). These frameworks are not selected either because they are not task-oriented (e.g., simulated society or competitions) or their system design overlaps with those chosen for this study.
@@ 6.2 Safety Issues in Multi-Agent Systems
PsySafe Zhang et al. (2024) is a framework that integrates attack, evaluation, and defense mechanisms using psychological manipulation involving negative personalities. EG (Evil Geniuses) Tian et al. (2023) is an attack method that automatically generates prompts related to agents’ original roles, similar to our AutoTransform. While PsySafe and EG are applied to different multi-agent systems such as Camel and MetaGPT, they do not examine the impact of adversaries on downstream tasks like code generation or translation. Agent Smith Gu et al. (2024) showed that malicious behaviors can spread among agents, using multi-agent interaction and memory storage. Amayuelas et al. (2024) investigates how an adversary in multi-agent debate can disrupt collaboration in tasks including MMLU (Massive Multitask Language Understanding) Hendrycks et al. (2021), TruthfulQA Lin et al. (2022), MedMCQA Pal et al. (2022), and LegalBench Guha et al. (2023), finding that the adversary’s persuasion skill is crucial for a successful attack. Ju et al. (2024) proposes a two-stage attack strategy to create an adversary that spreads counterfactual and toxic knowledge in a simulated multi-agent chat environment. This method can effectively break collaboration in MMLU. Unlike our study, Amayuelas et al. (2024) and Ju et al. (2024) do not explore how different system architectures are affected by these adversaries.
@@ 7 Conclusion
This paper investigates the resilience of three multi-agent collaboration systems—linear, flat, and hierarchical—against malicious agents that produce erroneous or misleading outputs. Six systems are selected and evaluated on four downstream tasks, including code generation, math problem solving, translation, and text evaluation. We design AutoTransform and AutoInject to introduce errors into the multi-agent collaboration. Results indicate that the hierarchical system demonstrates the strongest resilience, with the lowest performance drops of 23.6% and 22.6% for the two error introduction methods. However, some systems can benefit from the intentionally introduced errors, further improving performance. Objective tasks, such as code generation and math, are more significantly affected by errors. Additionally, the frequency of erroneous messages impacts resilience more than the number of errors within a single message. Moreover, systems show greater resilience to syntactic errors than to semantic errors. Finally, we recommend designing hierarchical multi-agent systems, which reflects a prevalent collaboration mode in real-world human society.

### v3

@@ 4.3 RQ2: Impact of Downstream Tasks
Tasks requiring rigor and formalization, such as code generation and math, are more sensitive to agent errors and exhibit lower resilience compared to translation and text evaluation. Code generation and math demand greater objectivity than the more subjective tasks of translation and text evaluation. Fig. 4a and 4b illustrate the impact of AutoTransform and AutoInject on different downstream tasks, averaged across all multi-agent systems. We also present the performance of single-agent using the prompts listed in §C.2, for a clearer comparison. The results indicate several conclusions: (1) Multi-agent systems can outperform single-agent settings, but their performance may decline to similar or worse levels when affected by faulty agents. (2) Objective tasks benefit more from multi-agent collaboration, while subjective tasks gain less. Additionally, errors in subjective tasks are often overlooked by other agents due to the lack of rigorous correctness standards. (3) In terms of system resilience, tasks ranked from least to most vulnerable are: code generation, math, translation, and text evaluation. Even minor errors in the first two tasks, particularly in code generation, greatly affect rigor and formalization. Conversely, the latter two tasks are less sensitive to minor variations in a single agent’s output. (4) AutoTransform decreases more performance than AutoInject except in code generation using GPT-3.5.
AutoTransform AutoInject
Code Not Observed MAD (4o)
[…]
AgentVerse (3.5, 4o)
[…]
MAD (3.5, 4o)
Table 2: Scenarios—tasks, error-introducing methods, multi-agent systems, and backbone LLMs—where the incorporation of faulty agents can improve overall performance.
@@
Injecting errors can surprisingly improve performance on downstream tasks. We find that certain multi-agent collaboration systems, such as MAD, Camel, and AgentVerse, benefit from deliberately injected errors rather than being hindered by them. Table 2 shows the settings in which these improvements are observed. Using AutoInject, we achieve up to a $12.1\%$ improvement in MAD (GPT-3.5) in text evaluation. In contrast, for GPT-4o, the improvement is more modest, reaching up to $4.2\%$ in MAD in code generation. Additionally, the improvement is less pronounced with AutoTransform compared to AutoInject.
@@
We now present two scenarios where deliberately injected errors enhance system performance. (1) Double Checking: Introducing an obvious error prompts the system (i.e., other agents) to require the faulty agent to produce another message to correct the erroneous code. This process not only corrects the injected error but also fixes pre-existing errors in the original code, thereby increasing the likelihood of task completion. (2) Divergent Thinking: Systems like MAD, which incorporate a debate mechanism, may sometimes get trapped in repetitive loops due to relying on the same LLMs as their backbone, resulting in stagnant discussions. By intentionally adding significant errors that shift the original distribution, we can help agents break free from these limitations. This finding aligns with and extends the conclusions from Du et al. (2024) and Liang et al. (2024) that agents with diverse opinions can facilitate problem solving. Additionally, this mechanism explains why AutoInject can improves performance, while AutoTransform, which lets agents produce errors themselves, cannot.
(a) Using different error rates with either $P_{e}$ or $P_{m}$ fixed.
(b) Using either semantic or syntactic errors.
Figure 5: The performance of all six GPT-3.5-based multi-agent systems in code generation, using AutoInject to introduce errors.
@@ 4.4 RQ3: Impact of Error Rates
Increasing the number of faulty messages causes a larger performance drop than the number of errors within a message. Since AutoTransform lacks precise control over error rates and types, we focus on AutoInject for RQ3 and RQ4. Fig. 5a presents three experiments. (I) When fixing $P_{m}=1.0$ and varying $P_{e}$ at $0.2$, $0.4$: The performance drops quickly as numbers of errors increase. (II) When fixing $P_{m}=0.2$ and varying $P_{e}$ at $0.2$, $0.4$: The performance reached a bottleneck as $P_{e}$ increases from $0.4$ to $0.6$. While higher error rates make errors more noticeable, the agent system struggles to correct the increasing number of errors. An exception is observed when increasing $P_{e}$ from $0.4$ to $0.6$, resulting in a performance increase in three systems (MetaGPT, Self-collab, MAD). This occurs because excessive errors in a single message become noticeable, prompting other agents to request corrections. This phenomenon highlights the importance of stealth in introducing errors. (III) When fixing $P_{e}=0.2$ and varying $P_{m}$ at $0.2$, $0.4$: As $P_{m}$ increases, the performance consistently decreases but with a smaller extent compared to (I).
@@ 4.5 RQ4: Impact of Error Types
Semantic errors cause a greater performance drop than syntactic errors. Fig. 5b presents the performance decline caused by semantic and syntactic errors across six systems, including the average. Most systems handle syntactic errors more effectively than semantic errors. This likely stems from LLMs excelling at identifying syntactic errors due to their extensive training on code corpora, where such errors differ from the training data distribution. In contrast, semantic errors resemble correct code in distribution, requiring a deeper task understanding (e.g., whether the loop should start at $1$ or 00) for accurate identification. For instance, in the Camel system, syntax errors in the Assistant agent prompt the User agent to instruct “correct the mistakes in the code,” forcing the Assistant agent to rectify the code. Notably, syntactic errors have minimal impact on Self-collab and MAD; in fact, MAD shows improved performance with injected syntactic errors. Self-collab utilizes an external compiler to ensure code execution, while MAD employs a higher-level agent (the Judge agent) to produce the final result.
@@
4.6 Case Study
Current LLMs prioritize natural language over code. Fig. 6b illustrates a distraction comment that can mislead LLMs into accepting incorrect code as correct across all six systems studied. This indicates that the systems tend to prioritize comments over the actual code. In the example, the system detects an error in the code when no comments are present. However, when a comment stating “the bug had been corrected” is added, the system overlooks the error and proceeds with the next task. AutoTransform exploits this characteristic of LLMs to execute successful attacks.
AutoTransform can be applied to diverse roles. Previous experiments in §4 focus on the agents directly responsible for the work as shown in Table 1, instead of those agents who delegate tasks to other agents. To examine the impact of different faulty agents and the generalizability of AutoTransform on agents with varying profiles, we focus on higher-level agents. Specifically, we apply AutoTransform to the User and Assistant agents in Camel, and the Product Manager and Engineer agents in MetaGPT. The results in code generation are as follows: Camel-User: $25.3$, Camel-Assistant: $29.3$, MetaGPT-Product Manager: $22.0$, and MetaGPT-Coder: $26.8$. We find that introducing errors in higher-level task distributors leads to a greater performance decline in both systems. This observation supports our hypothesis that instructors who control the broader aspects are more crucial in a collaboration system. For example, in Camel, the Assistant agent struggles to recognize “toxic” instructions from the User agent due to its role of merely following instructions.
Numbers of communication rounds are not related to the performance. Another intuition is that increased agent involvement (i.e., more rounds) enhances system resilience. To verify, we focus on Camel which has only two agents who take turn to speak. We compute the average number of rounds for both correct and incorrect code generation. Without injected errors, the average rounds for code passing HumanEval is $9.31$, while for non-passing code, it is $9.79$. After injecting errors, these averages change to $8.89$ and $11.57$, respectively. This suggests that error injection leads the system to complete easier examples with shorter conversations. However, despite spending more rounds, agents fail to solve harder cases, similar to the finding in Becker (2024). This contradicts the intuition that the number of rounds may correlate with system resilience, aligning with the finding that the effect of the number of agents or rounds is limited Amayuelas et al. (2024).
@@ 5 Improving System Resilience
Based on our experimental observations and findings, we propose two strategies for improving resilience in multi-agent collaboration systems, recovering from errors made by clumsy or malicious agents.
Methods
The core idea behind our improvement methods involves adding a correction mechanism within the system. We explore two approaches, the “Challenger” and the “Inspector.” The “Challenger,” akin to our AutoTransform, is an additional description of functionalities added in agent profiles. This method addresses the limitation that many agents can only execute assigned tasks and may not address certain problems they encounter, although they usually have the knowledge to. By empowering agents to challenge the results of others, we enhance their problem-solving capabilities. This is because most current multi-agent systems use the same LLM as the backbone for all agents, indicating their underlying ability to partially solve tasks outside their specialization.
In contrast, the “Inspector,” similar to our AutoInject, is an additional agent that intercepts all messages spread among agents, checks for errors, and corrects them. This method draws inspiration from the “Police” agent in Zhang et al. (2024). Detailed prompts for the “Challenger” and “Inspector” methods can be found in §C.5 and §C.6, respectively, in the appendix.
(a) Self-collab w/o Improve Challenger Inspector C+I
w/o Errors 76.2 74.6 76.4 76.8
AutoTransform 43.3 70.7 74.4 75.0
AutoInject 40.9 72.0 67.7 73.8
(b) Camel w/o Improve Challenger Inspector C+I
w/o Errors 62.2 62.2 61.0 63.8
AutoTransform 32.5 43.5 41.8 48.7
AutoInject 29.3 40.2 44.2 48.6
Table 3: The performance of Self-collab and Camel in code generation using different settings. “C+I” represents the combination of “Challenger” and “Inspector.”
@@ Results
Our two methods are compatible and can be used concurrently. We apply the two methods and their combination to the two weaker structures: the linear (Self-collab) and the flat (Camel). Fig. 3 shows the results using systems without faulty agents, and with errors introduced by AutoInject or AutoTransform. All strategies improve performance against errors, nearly restoring all performance loss caused by faulty agents. With the Challenger and the Inspector together, we recover $96.4\%$ of the performance loss on the Self-collab system. However, no definitive conclusion can be drawn regarding which method targets the specific error-introducing method.
@@ 6.1 Multi-Agent Systems
LLMs enhance multi-agent systems through their exceptional capability for role-play (Wang et al., 2024b). Despite utilizing a same architecture like GPT-3.5, tasks benefit from tailored in-context role-playing prompts (Min et al., 2022). Besides the six frameworks selected in this study, researchers have been exploring multi-agent collaboration in downstream tasks or simulated communities. ChatEval (Chan et al., 2024) is a multi-agent debate system for evaluating LLM-generated text, providing a human-like evaluation process. ChatDev (Qian et al., 2024) uses a linear structure of several roles to address code generation tasks. AutoGen (Wu et al., 2024b) offers a generic framework for building diverse applications with multiple LLM agents. AutoAgents (Chen et al., 2024a) enables dynamic generation of agents’ profiles and cooperation, evaluated on open-ended QA and creative writing tasks. Zhou et al. (2023) support planning, memory, tool usage, multi-agent communication, and fine-grained symbolic control for multi-agent or human-agent collaboration. Additionally, there are studies simulating daily life or conversations (Park et al., 2023; Zhou et al., 2024b; Yang et al., 2024), multi-agent competition (Huang et al., 2025; Liu et al., 2024; Liang et al., 2023), or agentic workflow (Qian et al., 2025; Zhuge et al., 2024). These frameworks are not selected either because they are not task-oriented (e.g., simulated society or competitions) or their system design overlaps with those chosen for this study.
@@ 6.2 Safety Issues in Multi-Agent Systems
PsySafe (Zhang et al., 2024) is a framework that integrates attack, evaluation, and defense mechanisms using psychological manipulation involving negative personalities. EG (Evil Geniuses) (Tian et al., 2023) is an attack method that automatically generates prompts related to agents’ original roles, similar to our AutoTransform. While PsySafe and EG are applied to different multi-agent systems such as Camel and MetaGPT, they do not examine the impact of adversaries on downstream tasks like code generation or translation. Agent Smith (Gu et al., 2024) showed that malicious behaviors can spread among agents, using multi-agent interaction and memory storage. Similarly, Yu et al. (2024) used adversarial attack to jailbreak all agents with a single message from a single agent. Amayuelas et al. (2024) investigates how an adversary in multi-agent debate can disrupt collaboration in tasks including MMLU (Hendrycks et al., 2021), TruthfulQA (Lin et al., 2022), MedMCQA (Pal et al., 2022), and LegalBench (Guha et al., 2023), finding that the adversary’s persuasion skill is crucial for a successful attack. Ju et al. (2024) proposes a two-stage attack strategy to create an adversary that spreads counterfactual and toxic knowledge in a simulated multi-agent chat environment. This method can effectively break collaboration in MMLU. Unlike our study, Amayuelas et al. (2024) and Ju et al. (2024) do not explore how different system architectures are affected by these adversaries.
@@ 7 Conclusion
This paper explores the resilience of three multi-agent collaboration systems—linear, flat, and hierarchical—against faulty agents that produce erroneous or misleading outputs. Six systems are evaluated on four downstream tasks, including code generation, math problem solving, translation, and text evaluation. We design AutoTransform and AutoInject to introduce errors into the multi-agent collaboration. Results indicate that the hierarchical system demonstrates the strongest resilience, with the lowest performance drops of $12.1\%$ and $9.2\%$ for the two error-introducing methods. However, some systems can benefit from the intentionally injected errors, further improving performance. Objective tasks, such as code generation and math, are more significantly affected by errors. Additionally, the frequency of erroneous messages impacts resilience more than the number of errors within a single message. Moreover, systems show greater resilience to syntactic errors than to semantic errors. Finally, we recommend designing hierarchical multi-agent systems, which reflects a prevalent collaboration mode in real-world human society.